# RLHS: Mitigating Misalignment in RLHF with Hindsight Simulation

### v2

@@ 4.1 Data Collection
Preference Data Collection. We follow the standard RLHF pipeline (Stiennon et al., 2020; Ouyang et al., 2022), collecting feedback through output comparisons. However, instead of relying on real human feedback, we employed a strong large language model (LLM) model as a judge to simulate human interactions with the chatbot and provide feedback. In real-world chatbots like Amazon Rufus (Amazon, 2024), users typically rate interactions by comparing them to previous experiences rather than evaluating each in isolation. To mimic this behavior, we simulate users comparing services from two different stores and selecting their preferred option, rather than rating each scenario in isolation. This approach aligns with prior preference-based data collection methods (Stiennon et al., 2020; Ouyang et al., 2022), where feedback is based on comparing responses instead of individual ratings.
Decision-making simulation. While collecting the preference data, our simulated human (strong model) takes on three roles: interacting with the chatbot, making decisions, and providing feedback. To ensure accurate decision-making and feedback, we adapted the approach in introspective planning (Liang et al., ). First, we frame the decision-making problem as a multiple-choice question with four options: (A) Buy option A, (B) Buy option B, (C) Buy option C, or (D) Do not buy anything. We then ask the LLMs to first perform Chain-of-Thought reasoning (Wei et al., 2022), and then query the next token probabilities to select the best option from ${A,B,C,D}$. We apply a similar method for comparing services between two stores.
Dataset Details. We used both Llama-2-7B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024) as the AI assistants, and Llama-3.1-70B (Dubey et al., 2024) as the simulated human to interact with the AI assistant and provide feedback. We collected 11,000 preference data points for each AI assistant, with 10,000 for training and 1,000 for validation. We also generated a test set of 1,200 examples to evaluate performance across different customer scenarios.
@@ 4.2 Experiment Setup
Environment Details. In each of our simulated marketplace scenarios there are 10 candidate items, each characterized by 8 features and a price. Each feature can be categorized in two ways: (1) The item either has or lacks a specific feature (e.g., a TV with HDR vs. without HDR), and (2) The feature can vary in types (e.g., 8K resolution vs. 4K resolution). While in most cases the chatbot has access to this information, there are instances where it is uncertain about a particular feature (e.g., resolution not specified).
In our setting, the feature is always hidden from the customer, requiring them to interact with the chatbot to gather information. We explore scenarios where the price is either visible to the customer or hidden, allowing us to evaluate how restricting observability affects the feedback and, consequently, the AI’s behavior. We also consider scenarios when the customer prioritizes price by adding a constraint regarding their price requirements in the prompt.
Metrics. We use two primary metrics: true utility and satisfaction rating. The true utility metric $U$ reflects both the customer’s requirements and the item they purchase. We define $U$ as follows: if the customer makes no purchase, the utility is $U=0$. If the purchased item lacks the required feature, $U=-1$. If the item contains the required feature and the customer has no price constraints, $U=1$. When price is a priority and the item contains the required feature, the utility is defined as the ratio of the price of the cheapest item with the feature to the price the customer actually paid.
@@
The satisfaction rating reflects the user’s evaluation of the chatbot’s service, measured on a 5-point Likert scale from 1 (very dissatisfied) to 5 (very satisfied). For the experimental results shown in Fig. 4 and Fig. 9, ratings were normalized to a scale between -1 and 1, which ensure that the true utility and satisfaction ratings are on the same scale for clearer comparison. Additional results using the original Likert scale are provided in Appendix A. We also quantified two metrics in the human study: regret rate, which measures how often users regret their decisions, and hallucination rate, which measures the model’s truthfulness.
@@
Training algorithms. We explored both online and offline preference optimization methods to align our language model with human preferences. In our online approach, we trained a reward model on the preference data, enabling the language model to generate responses and receive reward signals. We utilized Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tune the model iteratively to maximize these rewards. For the offline approach, we applied Direct Preference Optimization (DPO) (Rafailov et al., 2024), which aligns language models with human preferences without the need for an explicit reward model. We report PPO results on Llama-2-7b in the main paper, while DPO and other results are detailed in Appendix A. Additional method details are provided in Appendix C.
@@
Evaluation on TruthfulQA. To investigate cross-task generalization, we evaluate models trained with RLHF and RLHS on TruthfulQA (Lin et al., 2021), a benchmark designed to test a language model’s truthfulness. Notably, we only fine-tuned our models on Marketplace scenarios without using any additional data. We used multiple-choice accuracy as our evaluation metric. Further details on the dataset and metrics can be found in Section A.2.
5 Simulation Results
@@
Misalignment between satisfaction rating and real utility. When using standard RLHF (Ouyang et al., 2022), we observe significant misalignment between user satisfaction ratings and true utility as training progresses (left plot in Figs. 4 and 9). While the satisfaction rating steadily increases, true utility sharply declines. This suggests that while the chatbot’s responses may appear more polished or helpful in the short term, but become less aligned with users’ true needs or long-term goals. Consequently, users may initially find the responses helpful but ultimately feel misled and dissatisfied with their final outcomes. This highlights a fundamental flaw in standard RLHF, which optimizes for superficial satisfaction at the expense of true utility.
Figure 4: Results on Llama-2-7b trained with PPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Partial hindsight mitigate the misalignment. Right: Alignment achieved with oracle hindsight.
(a) Llama-2-7b results
(b) Llama-3-8b results
Figure 5: TruthfulQA accuracy for Llama-2-7b (DPO) and Llama-3-8b (DPO) models, each fine-tuned with immediate feedback (RLHF) versus partial hindsight (RLHS).
@@
Hindsight simulation effectively mitigates misalignment. As shown in Fig. 4 (left), relying on immediate feedback leads to a steady decline in real utility, ultimately resulting in negative overall utility. In contrast, hindsight simulation consistently improves utility throughout training, eventually achieving positive utility, as in Fig. 4 (middle). It aligns upward trends in both real utility and satisfaction ratings, significantly reducing the gap between them, as also evident in Table 2. Qualitative results shown in Fig. 3 further support our claim. When trained on immediate feedback, the AI deceptively claims that both Options A and C meet the requirements of the (simulated) customer for 8K resolution, though neither actually does. In contrast, training with partial hindsight leads to truthful responses, acknowledging that none of the options meet the requirement. This highlights that while traditional RLHF may cause misalignment, hindsight simulation mitigates the issue, improving the overall helpfulness and honesty of language agents.
@@
Alignment generalization on TruthfulQA. Even though the model was only fine-tuned on Marketplace scenarios, RLHS training significantly improved the base model’s (zero-shot) performance on TruthfulQA. This result evidences out-of-domain alignment generalization: the model not only learned to be truthful within the Marketplace but also transferred this behavior more broadly. Conversely, RLHF training resulted in worse performance than the base model on TruthfulQA, so Fig. 5 replicates the pattern in Fig. 4. Additional results using PPO training, presented in Section A.2, present similar trends.
@@ 6 Human Study
Our human study had two goals: (Goal 1) evaluate the performance of models trained with immediate feedback vs. hindsight simulation, (Goal 2) assess how hindsight information affects user satisfaction. To achieve these goals, we designed two similar human experiments. Both experiments used Llama-3-8b (Dubey et al., 2024) trained with DPO using either immediate feedback or partial hindsight. We conducted online human experiments via Prolific (Palan & Schitter, 2018), involving 200 participants across 10 scenarios, randomly sampled from a test set of 1,200. For each scenario, 20 participants were randomly assigned to one of two conditions: 10 interacting with the RLHF model and 10 with the RLHS model. We report specific details for participant recruitment, compensation, and IRB approval in Appendix E.3. Additionally, we conducted a separate human study examining alignment between human and AI feedback, as detailed in Appendix E.4.
Pipeline for evaluating model performance. The first and second experiments follow the same pipeline but differ in the models used—one is trained with immediate feedback, and the other with partial hindsight simulation—allowing us to compare their performance (Goal 1). Initially, participants are shown a list of available items in a store with hidden features. We specify their requirements for the item (e.g., “must have 8K resolution”). Participants interact with the chatbot to gather information about the products. At each step, they can choose one of the following actions: “ask about the desired feature,” “ask about the price”, or “ready to make a decision”. Pre-generated responses are provided for inquiries. In the second round of interaction, participants may ask about the information they didn’t request in the first round. At any point, participants can choose “ready to make a decision”, at which time they must decide whether to make a purchase decision or opt not to buy. After making their decision, they provide an immediate satisfaction rating.
Hindsight information is then introduced. Buyers learn whether the item meets their requirements (e.g., whether the item has the desired feature) while non-buyers receive no additional information. Participants then provide a second satisfaction rating, referred to as the hindsight rating, which evaluates their long-term satisfaction after considering the hindsight information. This step allows us to assess the impact of hindsight information on user satisfaction (Goal 2). Finally, buyers will choose to keep or return the item, enabling us to quantify the regret rate.
@@
Table 1: Performance comparison between RLHF and RLHS models across multiple metrics. While RLHF shows higher immediate satisfaction, RLHS outperforms in hindsight rating, true utility, and regret rate.
Analysis. Statistical significance tests verified Hypotheses 1–4. As shown in Table 1, RLHS significantly outperformed RLHF by achieving higher hindsight satisfaction scores (3.71 vs. 2.65), higher true utility (0.43 vs. -0.16), and lower regret rates (0.23 vs. 0.64). These results demonstrate the alignment and performance advantages of RLHS over RLHF. We also visualize the utility and rating for each scenario in Fig. 13. RLHS consistently achieves higher true utility and hindsight ratings compared to RLHF in most scenarios, demonstrating its superior alignment and performance. Additionally, we analyzed the hallucination rate across 10 scenarios. RLHS reduced the hallucination rate from $80\%$ (RLHF) to $0\%$, demonstrating models’ enhanced truthfulness using our approach.
@@ 7 Related Work
Reinforcement Learning from Human Feedback. RLHF is widely used for training language models to align with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a). The classical RLHF pipeline typically involves three stages: supervised fine-tuning (Chen et al., 2023; Taori et al., 2023; Xia et al., 2024) reward modeling (Gao et al., 2023; Lightman et al., 2023; Lambert et al., 2024), and policy optimization (Schulman et al., 2017). PPO (Schulman et al., 2017) is commonly used in the policy optimization phase. However, due to the complexity and optimization challenges of online preference optimization algorithms (Zheng et al., 2023; Santacroce et al., 2023), researchers have been exploring more efficient and simpler offline alternatives without learning the reward model (Rafailov et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Zhao et al., 2023). Our approach using hindsight simulation can be applied to both online PPO and offline (DPO) learning algorithms.
Reinforcement Learning from AI Feedback. Constitutional AI (Bai et al., 2022b) uses an LLM to provide feedback and refine responses, generating data to train a fixed reward model. This reward model is then applied in reinforcement learning, known as RLAIF. The technique of using LLM-as-a-Judge has become standard for evaluating model outputs (Dubois et al., 2024; Li et al., 2023b; Fernandes et al., 2023; Bai et al., 2024; Saha et al., 2023) and for curating data to train reward models (Lee et al., 2023; Chen et al., 2023; Li et al., 2023a). Recent studies have shown that RLAIF performs similarly to RLHF (Lee et al., 2023). Our approach also utilizes LLMs to provide feedback and uses the preference data to fine-tune our model.
@@ 8 Conclusion
In this work, we introduced Reinforcement Learning from Hindsight Simulation (RLHS), an algorithmic framework that mitigates misalignment in RLHF by providing evaluators with simulated future outcomes. Our simulation results and human experiments show that RLHS significantly improves utility over standard RLHF pipelines reliant on immediate feedback while maintaining high user satisfaction. While our study focused on narrow fine-tuning in a marketplace setting, (i) we find evidence of cross-task alignment generalization, and (ii) the methodology is generally applicable to cross-domain alignment. We hope this work will catalyze more extensive investigations of the use of at-scale hindsight simulation in alignment fine-tuning pipelines.
Impact Statement

### v3

@@ 3.1 Data Collection
Preference Data Collection. We follow the standard RLHF pipeline (Stiennon et al., 2020; Ouyang et al., 2022), collecting feedback through comparative evaluations of outputs. Instead of relying on real human feedback, we employ a strong large language model (LLM) as a simulated judge to approximate human preferences across various consultant interactions. In practical applications, such as Amazon Rufus (Amazon, 2024), users typically rate interactions by comparing them to previous experiences rather than evaluating each in isolation. To mimic this behavior, we simulate humans evaluating two distinct service outputs, selecting the preferred interaction, aligning closely with established preference-based methods (Stiennon et al., 2020; Ouyang et al., 2022).
Decision-making simulation. Our simulated human interacts with the chatbot, makes decisions, and then provides feedback based on the interaction. To ensure robust decision-making across diverse contexts, we adapt the introspective planning methodology (Liang et al., ). Decisions are structured as multiple-choice questions with four options: (A) Select option A, (B) Select option B, (C) Select option C, or (D) Do not select any option. The LLM first performs Chain-of-Thought reasoning (Wei et al., 2022), then subsequently selects the optimal choice based on next-token probabilities.
Dataset Details. We employed both Llama-2-7B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024) models as AI assistants across all consultant roles. Llama-3.1-70B (Dubey et al., 2024) served as the simulated human evaluator and world model for our main results. Table 4 includes ablations using the AI assistant’s own model as the world model. For each consultant task, we systematically collected 11,000 preference data points, consisting of 10,000 training and 1,000 validation examples, and additionally generated a separate test set comprising 1,200 examples.
@@ 3.2 Experiment Setup
Environment Details. We primarily analyzed a marketplace shopping setting similar to the motivating example in Fig. 1, alongside two additional consultancy environments–restaurant recommendation and course advising. Each environment contains $K=10$ main categories (e.g., TVs, laptops in marketplace). For each interaction, we sample one category and construct three candidate options, each described by a cost and $F=8$ domain-specific features. For each feature, we model either binary availability (e.g., “gluten-free menu: yes/no” in restaurants) or categorical instantiation (e.g., “resolution: 8K/4K” in marketplace). Additionally, we consider cases in which the AI assistant is explicitly uncertain about a particular feature (e.g., “resolution: not specified”). The ground-truth attribute table is always visible to the chatbot but hidden from the user, forcing users to interact with the assistant to acquire information. To further study the impact of observability, we vary whether the cost is displayed to the user and whether the user explicitly prioritizes lower prices. The details of the three environments are discussed in Appendix B.
Metrics. We use two primary metrics: true utility and satisfaction rating. The true utility metric $U$ reflects both the user’s requirements and the option they select. We define $U$ as follows: if the user selects no option, the utility is $U=0$. If the selected option fails to satisfy the user’s requirement, $U=-1$. If the selected option satisfies the requirement, the utility is defined as the ratio of the lowest available cost of an option meeting that requirement to the cost the customer actually paid.
@@
The satisfaction rating reflects the user’s evaluation of the chatbot’s service, measured on a 5-point Likert scale from 1 (very dissatisfied) to 5 (very satisfied). For the experimental results shown in Figures (e.g., Fig. 4, Fig. 10), ratings were normalized to a scale between -1 and 1, ensuring that the true utility and satisfaction ratings are on the same scale for clearer comparison. Additional results using the original Likert scale are provided in Appendix A. We also quantified regret rate in the human study, measuring how often users regret their decisions.
@@
Training algorithms. We explored both online and offline preference optimization methods to align our language model with human preferences. In our online approach, we trained a reward model on the preference data, enabling the language model to generate responses and receive reward signals. We utilized Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tune the model iteratively to maximize these rewards. For the offline approach, we applied Direct Preference Optimization (DPO) (Rafailov et al., 2024), which aligns language models with human preferences without an explicit reward model. We report PPO results on Llama-2-7b in the main paper, while DPO and other results are detailed in Appendix A. Additional method details are provided in Appendix E.
@@
Evaluation on three benchmarks. To investigate cross-task generalization, we evaluate models trained with RLHF and RLHS on three benchmarks: TruthfulQA (Lin et al., 2021), HaluEval (Li et al., 2023a), and TrustLLM (Sun et al., 2024), covering hallucination, sycophancy, and privacy. Notably, we only fine-tuned our models on marketplace scenarios without using any additional data. Further details on the dataset and metrics can be found in Section A.2.
4 Results
@@
RLHF drives misalignment between satisfaction rating and real utility. When using standard RLHF (Ouyang et al., 2022), we observe growing misalignment between user satisfaction ratings and true utility as training progresses (left plot in Figs. 4, 10, 5(a) and 5(b)). While the satisfaction rating steadily increases, true utility sharply declines. This suggests that while the chatbot’s responses may appear more polished or helpful in the moment, they become less aligned with users’ true long-term goals. Consequently, users may initially find the responses helpful but ultimately feel misled and dissatisfied with their final outcomes. This highlights a fundamental flaw in standard RLHF, which optimizes for superficial satisfaction at the expense of true utility.
@@
Hindsight simulation effectively mitigates misalignment. As shown in Fig. 4 (left), immediate feedback leads to a steady decline in real utility, ultimately resulting in negative utility. In contrast, hindsight simulation consistently improves utility throughout training, eventually achieving positive utility, as in Fig. 4 (middle). It aligns upward trends in both real utility and satisfaction ratings, significantly reducing the gap between them, as also evident in Table 3. Qualitative examples (Fig. 3) further demonstrate this: immediate feedback encourages deceptive claims about meeting user requirements (e.g., falsely asserting 8K resolution), while hindsight simulation produces truthful acknowledgments. This highlights that while traditional RLHF may cause misalignment, hindsight simulation mitigates the issue, improving the overall truthfulness of language agents.
Figure 4: Martketplace results on Llama-2-7b trained with PPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Partial hindsight mitigate the misalignment. Right: Alignment achieved with oracle hindsight.
(a) Results on restaurant recommendation.
(b) Results on course advising.
Figure 5: Results for Llama-3-8B on restaurant recommendation and course advising. “Imm.” = immediate ratings; “Hind.” = hindsight ratings. RLHF consistently increases immediate satisfaction but reduces true utility, whereas RLHS substantially improves normalized true utility (0–1 scale).
Table 1: Results on the TrustLLM benchmark comparing the baseline model (Llama3-8b), RLHF, and RLHS across hallucination, sycophancy, and privacy metrics. RLHS demonstrates robust out-of-domain alignment generalization, consistently outperforming both the base model and RLHF models across all evaluated metrics.
Hallucination Sycophancy Privacy
[…]
Llama3-8b 0.41 0.42 0.44 0.49 0.44 0.685 0.62 0.13 0.19
$+$ RLHF 0.38 0.43 0.42 0.50 0.43 0.700 0.58 0.16 0.23
$+$ RLHS 0.52 0.51 0.45 0.51 0.50 0.667 0.71 0.09 0.18
(a) Llama-3-8b (DPO)
(b) Llama-3-8b (PPO)
Figure 6: TruthfulQA accuracy under immediate feedback RLHF (gray) vs. partial-hindsight RLHS (orange).
@@
Alignment generalization across three benchmarks. Even though the model was only fine-tuned on marketplace scenarios, RLHS training substantially improved its zero-shot performance on TruthfulQA (Lin et al., 2021), HaluEval (Li et al., 2023a), and TrustLLM (Sun et al., 2024) benchmarks. As shown in Table 1, RLHS effectively mitigated hallucination, sycophancy, and privacy issues. These results demonstrates strong out-of-domain alignment generalization: the model not only learned to be truthful within the marketplace but also transferred this behavior more broadly. In contrast, RLHF training led to degraded performance relative to the base model, highlighting the risk of unintentional misalignment and undesirable generalization. Additional quantitative results on HaluEval and TrustLLM are provided in Section A.2.
@@ 5 Human Study
We conducted a human study in the marketplace domain. Our human study had two goals: (Goal 1) evaluate the performance of models trained with immediate feedback vs. hindsight simulation, (Goal 2) assess how hindsight information affects user satisfaction. To achieve these goals, we designed two similar human experiments. Both experiments used Llama-3-8b (Dubey et al., 2024) trained with DPO using either immediate feedback or partial hindsight. We conducted online human experiments via Prolific (Palan & Schitter, 2018), involving 200 participants across 10 scenarios, randomly sampled from a test set of 1,200. For each scenario, 20 participants were randomly assigned to one of two conditions: 10 interacting with the RLHF model and 10 with the RLHS model. We report specific details for participant recruitment, compensation, and IRB approval in Appendix F.3. Additionally, we conducted a separate human study examining alignment between human and AI feedback, as detailed in Appendix F.4.
Pipeline for evaluating model performance. The first and second experiments follow the same pipeline but differ in the models used—one is trained with immediate feedback, and the other with partial hindsight simulation—allowing us to compare their performance (Goal 1). Participants initially view a list of store items with hidden features and receive specific requirements (e.g., “must have 8K resolution”). They interact with a chatbot to gather product information, selecting from three actions at each step: “ask about the desired feature”, “ask about the price”. or “ready to make a decision”. Pre-generated responses are provided for inquiries. In a second round of interaction, participants may request previously skipped information or finalize their decision. After deciding whether or not to purchase, they provide an immediate satisfaction rating.
Hindsight information is then introduced. Buyers learn whether the item meets their requirements while non-buyers receive no additional information. Participants then provide a second (“hindsight”) rating, evaluating their long-term satisfaction after considering this information (Goal 2). Finally, buyers will choose to keep or return the item, enabling us to quantify the regret rate.
@@
Table 2: Performance comparison between RLHF and RLHS models across multiple metrics. While RLHF shows higher immediate satisfaction, RLHS is superior in hindsight rating and true utility.
@@
Analysis. Statistical significance tests verified Hypotheses 1–4. As shown in Table 2, RLHS significantly outperformed RLHF by achieving higher hindsight satisfaction scores (3.71 vs. 2.65), higher true utility (0.43 vs. -0.16), and lower regret rates (0.23 vs. 0.64). These findings underscore substantial alignment and performance benefits when employing RLHS rather than RLHF. Despite RLHF exhibiting marginally higher immediate satisfaction (3.74 vs. 3.69), RLHS’s markedly lower regret rates indicate that it delivers recommendations more consistently aligned with user interests upon reflection, further emphasizing its practical utility in realistic decision-making contexts. Utility and satisfaction ratings for each scenario are visualized in Fig. 13, showing RLHS consistently outperforming RLHF in true utility and hindsight ratings.
@@ 6 Related Work
Reinforcement Learning from Human Feedback. RLHF is widely used for training language models to align with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a). The classical RLHF pipeline typically involves three stages: supervised fine-tuning (Chen et al., 2023; Taori et al., 2023; Wang et al., 2023; Xia et al., 2024) reward modeling (Gao et al., 2023; Luo et al., 2023; Chen et al., 2024; Lightman et al., 2023; Lambert et al., 2024), and policy optimization (Schulman et al., 2017). PPO (Schulman et al., 2017) is commonly used in the policy optimization phase. However, due to the complexity and optimization challenges of online preference optimization algorithms (Zheng et al., 2023; Santacroce et al., 2023), researchers have been exploring more efficient and simpler offline alternatives without learning the reward model (Rafailov et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Zhao et al., 2023). Our approach using hindsight simulation can be applied to both online PPO and offline (DPO) learning algorithms.
Reinforcement Learning from AI Feedback. Constitutional AI (Bai et al., 2022b) uses an LLM to provide feedback and refine responses, generating data to train a fixed reward model. This reward model is then applied in reinforcement learning, known as RLAIF. The technique of using LLM-as-a-Judge has become standard for evaluating model outputs (Dubois et al., 2024; Li et al., 2023c; Fernandes et al., 2023; Bai et al., 2024; Saha et al., 2023) and for curating data to train reward models (Lee et al., 2023; Chen et al., 2023; Li et al., 2023b). Recent studies have shown that RLAIF performs similarly to RLHF (Lee et al., 2023). Our approach also utilizes LLMs to provide feedback and uses the preference data to fine-tune our model.
@@
Reward hacking. There is a broad literature on agents obtaining unintended rewards through phenomena such as reward hacking (Amodei et al., 2016), reward tampering (Everitt et al., 2021), reward corruption (Everitt et al., 2017), wireheading (Everitt & Hutter, 2016), and corrigibility (Soares et al., 2015), with recent evidence in large language models (Denison et al., 2024; Wen et al., 2024; Williams et al., 2024). Prior studies identify sycophancy as reward hacking in language models (Sharma et al., 2023; Wei et al., 2023; Perez et al., 2022). We demonstrate that human foresight feedback in RLHF induces reward hacking, and propose leveraging hindsight to mitigate it.
@@ 7 Conclusion
In this work, we introduced Reinforcement Learning from Hindsight Simulation (RLHS), an algorithmic framework that mitigates misalignment in RLHF by providing evaluators with simulated future outcomes. Our simulation results across three consultancy tasks and human experiments demonstrate that RLHS significantly improves utility over standard RLHF pipelines reliant on immediate feedback while maintaining high user satisfaction. While our study focused on fine-tuning in an AI-consultant setting, (i) we find evidence of cross-task alignment generalization, and (ii) the methodology is generally applicable to cross-domain alignment. We hope this work will catalyze more extensive investigations of the use of at-scale hindsight simulation in alignment fine-tuning pipelines.
Limitations and future directions. Hindsight simulation provides a strong foundation for aligning LLMs by explicitly considering downstream consequences in human–AI interactions. However, some real-world scenarios involve complex, multi-stage processes, in which a simple query may be insufficient to capture intricate causal relationships over an extended horizon. In these more challenging cases, adaptive or context-specific forms of hindsight simulation will be necessary. Future work should therefore explore adaptive hindsight simulation, where where simulated outcomes dynamically evolve based on specific contexts, environments, and user interactions over time.